OpenStudy (anonymous):

Graph: 1/2x^2 +5 ! who ever helps me with this will surely be blessed:)

OpenStudy (anonymous):

so graph it on the calculator

OpenStudy (anonymous):

i would but i mean she wants to see it on a paper.

OpenStudy (anonymous):

graph it then draw it

OpenStudy (anonymous):

but still on the test if im not given a graphing calculator how am i going to figure it out?

OpenStudy (anonymous):

every test lets you have a calculator period.

OpenStudy (anonymous):

not a graphing one.

OpenStudy (anonymous):

plug in x values and y values to get points

OpenStudy (anonymous):

how do i get my vertex?

OpenStudy (anonymous):

Find the vertex of y = 3x2 + 2x – 1

OpenStudy (anonymous):

Your original equation does not have a vertex: http://www.wolframalpha.com/input/?i=y%3D1%2F2x%5E2%2B5

OpenStudy (anonymous):

thank you jeremy! :D

OpenStudy (helder_edwin):

do u have to graph \[ \large y=\frac{1}{2}x^2+5 \]

OpenStudy (helder_edwin):

well if it's so u can rewrite this as \[ \large y-5=\frac{1}{2}x^2 \] it is a parabola (only one variable squared) with vertex (0,5)

OpenStudy (helder_edwin):



OpenStudy (anonymous):

thanks for the effort! <3

OpenStudy (helder_edwin):

u r welcome. don't hate your math teachers :)

OpenStudy (anonymous):

i wont. lol :)

OpenStudy (helder_edwin):

great :) study hard. always.